OpenStudy (anonymous):

A 5.5 kg bowling ball initially at rest is dropped from the top of a 12 m building. It hits the ground 1.75 s later. Find the net external force on the falling ball.

OpenStudy (yrelhan4):

s=1/2at^2 you have s and t find a. net force=ma. plug and solve.

OpenStudy (anonymous):

What is s?

OpenStudy (yrelhan4):

s-->distance travelled a--> acceleration due to gravity here t-->time taken

OpenStudy (anonymous):

Okay, thanks:)

OpenStudy (yrelhan4):

cheers!

OpenStudy (anonymous):

I 431.2N but that isn't an option..

OpenStudy (yrelhan4):

43.12 N is the answer.

OpenStudy (anonymous):

So this is what I did.. s=12m a=? t=1.75s 12=((1.75)^2a)/2 12=(3.0625a)/2 24=3.0625a a=7.84 netforce=55*7.84 netforce=431.2N

OpenStudy (yrelhan4):

Net acceleration. a = 2*12/1.75² Net force ma = 5.5* 2*12/1.75² = 43.12 N

OpenStudy (anonymous):

I get what your saying now. But the only problem is my choices are 120N, 54N, & 0N. But I can rule out 0N automatically

OpenStudy (yrelhan4):

hmm. no idea. sorry.

OpenStudy (anonymous):

Its okay, thanks for your help though :) I really do appreciate it

OpenStudy (yrelhan4):

you're welcome. good luck. :)